OpenStudy (anonymous):

If f(x)=x to the 2nd power minus 3 and f(x)=46,what is the value of x?

OpenStudy (zzr0ck3r):

\(x^2-3=46\implies x^2=49\implies x=\pm\sqrt{49}=\pm \ ? \)

OpenStudy (solomonzelman):

±7